# Adaptation of Agentic AI

### v2

@@ 5 Tool Adaptation
Tool adaptation represents another emerging trend for enhancing AI agents’ performance on specific tasks, marking a conceptual shift from optimizing the agent itself to optimizing its ecosystem. Instead of modifying the agent’s parameters through fine-tuning or reinforcement learning, this paradigm targets the external components—the “tools” that mediate perception, computation, and interaction. These tools may encompass pre-trained models, retrievers, planners, or executors that the agent can invoke through language or code. Consequently, tool adaptation focuses on improving the agent’s operational environment rather than its internal cognition. Methods in this category typically (1) employ pre-trained machine learning models—developed via environment feedback or data imitation—as plug-and-play components, ranging from simple classifiers to complex LLM-based sub-agents as discussed in §4; or (2) leverage the agent’s outputs as supervision or reinforcement signals to train, align, or refine the tool itself. This perspective re-frames intelligent systems as co-adaptive ecosystems, where frozen agents and adaptive tools evolve symbiotically toward higher task efficiency, modularity, and generalization.
@@
where $\mathcal{O}_{\text{tool}}$ quantifies task-specific or environment-driven improvements that are independent of the agent, such as retrieval accuracy or planning efficiency, while $\mathcal{O}_{\text{agent}}$ incorporates agent-derived supervision, where the agent’s outputs provide learning signals to refine or align the tool. Here, $\mathcal{T}^{*}$ denotes the optimized tool configuration that maximizes the respective objective, illustrating how tool adaptation complements agent-level optimization within the broader agent–tool ecosystem.
@@ 5.1 T1: Agent-Agnostic Tool Adaptation
The foundational architecture for tool-augmented systems is to use pre-trained models as plug-and-play tools for frozen agents. The agent orchestrates tool usage through prompting alone, never updating its parameters, while leveraging tools that were trained independently on diverse data sources before deployment.
@@ 5.1.1 Foundational Systems and Architectures
Early systems exemplifying the tool-adaptation paradigm established the architectural foundations for how frozen agents can effectively orchestrate or invoke external models. These pioneering works demonstrate distinct mechanisms—functional, prompt-based, code-based, and graph-based—that together shaped the design space for modern tool-augmented AI systems.
@@
Operator-Learning Tools: Neural Operators [124] (JMLR). Before large-scale LLM-based orchestration emerged, Neural Operators represented a seminal example of agent-agnostic tool learning: models trained to approximate mappings between infinite-dimensional function spaces, serving as differentiable surrogates for complex simulators. Unlike conventional neural networks tied to discrete grids, Neural Operators are discretization-invariant—they learn the underlying operator itself, not its finite discretization—and can generalize across resolutions and geometries. The Fourier Neural Operator (FNO) achieves $\mathcal{O}(J\log J)$ inference via spectral convolution and outperforms classical solvers on Navier–Stokes, Darcy flow, and elasticity equations by orders of magnitude in speed. Conceptually, FNO and its variants (Graph-, Low-rank-, Multipole-NO) mark the first wave of “frozen tools” that agents can query repeatedly for reasoning, planning, or control without retraining. In modern agentic pipelines, they are used as plug-in surrogates—fast, differentiable black-box functions invoked within decision or inference loops.
@@
HuggingGPT [125] (NeurIPS 2023) pioneered the orchestration paradigm by enabling ChatGPT to command 1000+ machine learning models from HuggingFace Hub without any fine-tuning. The frozen LLM executes a four-stage workflow: task planning (decomposing user requests), model selection (choosing from tool descriptions), task execution (invoking models), and response generation (synthesizing outputs). This architecture demonstrates that language serves as a universal interface—tool descriptions in natural language suffice for the frozen agent to coordinate complex multimodal workflows. On composite cross-modal tasks, HuggingGPT enables GPT-3.5 to achieve performance comparable to GPT-4V by orchestrating specialized vision, speech, and language models. The primary limitation lies in latency from sequential LLM calls and token length constraints for tool descriptions. ViperGPT [126] (ICCV 2023) introduced code generation as the orchestration mechanism. The frozen GPT-3 Codex generates Python code that composes vision models—GLIP for detection, SAM for segmentation, MiDaS for depth estimation—into executable programs. This code-based approach achieves state-of-the-art zero-shot performance on GQA visual reasoning, outperforming end-to-end models by 10–15% on compositional tasks. The key insight: Python functions provide more flexible tool composition than fixed API calls. Each tool exposes simple functions like find(image, object_name) or compute_depth(image), which Codex chains programmatically without learning tool-specific interfaces. SciToolAgent [127] (Nature Computational Science 2025) scales tool orchestration to scientific domains through graph-based organization. The frozen GPT-4o accesses 500+ biology, chemistry, and materials science tools via SciToolKG—a knowledge graph encoding tool metadata, dependencies, and safety constraints. Graph-based retrieval for tool selection achieves 94% accuracy on scientific query benchmarks, representing a 15–20% improvement over GPT-4o without tool access. The system successfully automates protein engineering workflows chaining ESMFold for structure prediction, BLAST for sequence alignment, and custom analysis tools. This architecture demonstrates that structured knowledge graphs address scalability challenges inherent in prompt-based descriptions.
@@
These foundational systems illustrate the dominant integration patterns. HuggingGPT exemplifies prompt-based orchestration, where the agent parses tool calls from text. ViperGPT uses code generation, exposing tools as Python functions. SciToolAgent demonstrates knowledge graph retrieval, using RAG to select from structured tool graphs. A fourth common pattern is multimodal bridging, which converts non-textual modalities into text representations; for example, Visual ChatGPT’s [128] prompt manager serializes vision operations as text API calls. The usability of these patterns depends on clear interface design, such as programmatic function signatures (e.g., find_object(image: PIL.Image, ...)), structured JSON schemas, or simple natural language descriptions. Execution modes are similarly varied, ranging from direct API calls and code generation to HTTP requests and command-line invocations.
@@
Model Context Protocol (MCP) and Code Execution Environments [129]. As large-scale agent ecosystems began to connect thousands of heterogeneous tools, the Model Context Protocol (MCP) emerged as an open standard for unifying how agents interface with external systems. Rather than embedding long tool definitions and intermediate results directly into the model’s context, MCP provides a universal API layer that enables frozen agents to discover, invoke, and coordinate tools across domains using a consistent schema. Building upon this infrastructure, Anthropic’s “Code Execution with MCP” paradigm introduced an execution-centric design in which the agent writes executable code to interact with MCP servers instead of performing token-level tool calls. This approach allows agents to load only the necessary tool definitions, filter or aggregate data within a sandboxed environment, and pass compact results back to the model, reducing context usage by over 98% while maintaining full compositionality. Conceptually, MCP represents a scalable T1-style tool adaptation infrastructure that decouples execution from inference, while the code-execution mode bridges toward T2-style optimization by dynamically improving efficiency under frozen agents. Together with systems such as HuggingGPT, ViperGPT, and SciToolAgent, it exemplifies the architectural evolution from static tool invocation to protocol-driven, programmable, and context-efficient orchestration.
@@ 5.1.2 Categories and Training Methodologies
Tool adaptation encompasses a wide range of pre-trained model categories, each contributing unique functional capabilities to frozen agents. The following examples highlight representative tool families and the training methodologies that enable their plug-and-play deployment across multimodal and domain-specific settings.
@@
Vision models dominate T1 deployments as plug-and-play tools. CLIP [130], trained contrastively on 400M image-text pairs, provides zero-shot classification and semantic understanding through frozen encoders. SAM [131], trained on 11M images with 1B masks via human-in-the-loop data engines, enables promptable segmentation with point, box, or mask inputs. SAM-CLIP [132] merges these capabilities through multi-task distillation with frozen teachers, achieving +6.8% mIoU on Pascal VOC for zero-shot semantic segmentation while retaining both parent models’ strengths. These vision tools require no task-specific fine-tuning - frozen agents invoke them directly via APIs for image understanding, segmentation, and classification tasks.
@@
Speech and audio tools leverage massive pre-training for robust performance. Whisper [133], trained on 680K hours of multilingual audio with weak supervision, provides speech recognition, translation, and language identification as a frozen API for multimodal agents. The encoder-decoder Transformer architecture enables zero-shot transcription across languages and domains, demonstrating remarkable robustness to accents, noise, and technical terminology. Agents simply pass audio inputs to the frozen Whisper model and process text outputs without any model adaptation.
@@
Code execution tools encompass models that learn to compose and execute functions through code. CodeAct [57] demonstrates that representing tool use in executable Python rather than static JSON improves compositional reasoning, achieving over 20% higher success rates on API-Bank benchmarks. The dynamic nature of code allows agents to flexibly construct, parameterize, and combine tools without predefined schemas.
@@
Search and retrieval tools comprise pre-trained dense retrievers such as DPR [134], ColBERT [135], Contriever [136], and e5 [137], often deployed as frozen components within retrieval-augmented generation pipelines. These bi-encoder models, trained on passage ranking tasks, enable semantic search over large corpora.
@@
Scientific tools extend capabilities to specialized fields. AlphaFold2 [138] and ESMFold [139] provide protein structure prediction from sequences. Materials science models like CGCNN [140] predict crystal properties. Molecule representation learning approaches [141, 142, 143, 144, 145, 146, 147] have been developed for molecular property prediction, whereas some encoder–decoder frameworks [148, 149] aim to predict transcriptional profiles elicited by chemical perturbations. These tools represent years of domain-specific model development, deployed as-is for frozen agents tackling scientific queries.
@@ 5.2 T2: Agent-Supervised Tool Adaptation
The T2 paradigm represents a profound conceptual inversion in how we approach adaptation in agentic systems. Rather than asking “how can we modify the agent to better use its tools?” (the A1/A2 question), T2 asks: “how can we modify the tools to better serve a fixed agent?” This seemingly simple reversal has far-reaching implications. It reframes the expensive, monolithic foundation model as a stable source of supervision rather than the target of optimization, and reconceptualizes the agent’s ecosystem of tools as a dynamic, adaptive periphery that can be continuously refined. This inversion is not merely a technical choice but reflects a deeper understanding of the economics and modularity of modern AI systems. Training or fine-tuning billion-parameter foundation models is computationally prohibitive and risks catastrophic forgetting. In contrast, the peripheral tools—retrievers, planners, memory modules—are typically orders of magnitude smaller and can be trained with dramatically less data and computation. The T2 paradigm exploits this asymmetry, achieving what we term symbiotic adaptation: the frozen agent provides high-quality supervision signals derived from its vast pre-trained knowledge, while the tools learn to translate, filter, and present information in exactly the form the agent finds most useful.
The evolution of T2 methods from 2023 to 2025 reveals a clear intellectual progression: from using internal proxy signals (perplexity, preferences) to train passive retrieval tools, to using verifiable outcome signals (task success, accuracy gains) to train active, multi-turn agentic tools. This progression mirrors a broader maturation in the field’s understanding of what makes an effective training signal and what kinds of tools are worth building.
@@ 5.2.1 Earlier Methods: From Proxy Signals to Structured Preferences
The earliest T2 methods emerged from the retrieval-augmented generation (RAG) community, where researchers sought to optimize dense retrievers for compatibility with large language models. These foundational works established the core principle that a frozen LM’s internal computations could serve as supervision, but they reveal, in hindsight, the limitations of relying on proxy metrics that may not align with downstream task objectives.
REPLUG [11] (NAACL 2024) introduced a general framework for adapting frozen language models through black-box supervision, using perplexity reduction as a training signal for the retriever. The core intuition is that if conditioning the LM on a retrieved document lowers its perplexity for a given query, the document likely provides informative context. Formally, the retriever is optimized to align its retrieval distribution with the distribution induced by the LM’s perplexity-based preferences:
@@
where $P_{\text{LM-perplexity}}(d|q)$ reflects how strongly each document reduces the LM’s perplexity when conditioned on query $q$. This design enabled retrieval adaptation without parameter access to the LM, establishing the foundation for a family of agent-supervised methods that optimize external modules based solely on frozen-agent feedback. BLADE [150] (AAAI 2025) further extended this paradigm by replacing traditional retrievers with domain-specific models that synthesize auxiliary knowledge: it couples a frozen general LLM with a small, domain-specific LM optimized via Bayesian Prompted Optimization (BPO). The small LM learns to generate domain-relevant knowledge and soft prompts that improve the black-box LLM’s responses, extending REPLUG’s black-box adaptation principle from retrieval to generative, domain-specialized tool co-adaptation. BBox-Adapter [151] (ICML 2024) reframed adaptation as an energy-based modeling problem, introducing a ranking-based noise-contrastive estimation loss and an online update framework to align outputs from black-box APIs like GPT-3.5 without access to internal probabilities. These methods collectively shifted the focus from likelihood-based alignment to utility-driven adaptation, paving the way for reinforcement-learning-based search and reasoning frameworks. proxy-tuning [152] (COLM 2024) extends black-box, agent-supervised adaptation to decoding time: a small tuned “expert” and its untuned “anti-expert” provide logit offsets that steer a frozen large LM without modifying its weights, effectively training a lightweight steering tool under a frozen agent (T2). EVOR [153] (EMNLP 2024) further extends to the domain of code generation, formulating retrieval and knowledge evolution as a co-adaptive process driven by execution feedback from a frozen LLM.
@@
While black-box adaptation methods such as REPLUG and BBox-Adapter relied on indirect proxy signals like perplexity or ranking scores, subsequent studies [28, 154, 155, 156, 157] moved toward explicit preference-based supervision that better reflects task utility. AAR [28] (ACL 2023) introduced augmentation-aware retrieval, where a frozen source LM (Flan-T5-250M) constructs preference pairs by comparing documents that most improve its own likelihood against human-annotated references. The retriever is then trained to reproduce these preferences via a contrastive loss, yielding a signal that directly encodes the LM’s notion of “helpful context.” Remarkably, these LM-derived preferences transfer effectively across architectures and scales, improving even 175B-parameter models. RA-DIT [154] (ICLR 2023) formalized this idea by defining document utility as the log-probability gain for producing correct answers:
@@
training retrievers to prefer documents that yield higher expected gains. This preference-based approach aligns retrieval more directly with downstream reasoning objectives while still operating through a single forward pass of the frozen LM. Together, these works mark a conceptual shift from proxy-based to task-aligned supervision, paving the way for reinforcement-style feedback and multi-turn optimization in later frameworks.
Multi-Stage Architectures: Distilling Complex Preferences
@@
The sophistication of training signals reached a new level with LLM-R [23] (EACL 2024), which introduced a multi-stage distillation pipeline. Rather than training the retriever directly on the frozen LM’s outputs, LLM-R first trains an intermediate cross-encoder “reward model” to capture the frozen LM’s nuanced preferences over in-context examples. This reward model, which can afford to be slow and expressive because it’s only used during training, is then distilled into a fast bi-encoder retriever. This architecture embodies a key design principle: the complexity of the training signal need not be constrained by inference-time efficiency requirements. By decoupling the preference modeling from the final retrieval tool, LLM-R achieves both high-quality supervision and fast deployment. UPRISE [158] (EMNLP 2023) extends this paradigm beyond documents to prompts, training a prompt retriever using the frozen LLM’s task performance across diverse tasks. By training on multiple tasks simultaneously, UPRISE learns a generalizable meta-skill: selecting prompts that improve LLM performance in zero-shot settings. The cross-task transfer (+8.5% on reading comprehension, +14.6% on paraphrase detection) suggests that T2-trained tools can internalize abstract principles of “what helps an LM” rather than memorizing task-specific heuristics.
@@
By 2024, the field had reached a critical realization: optimizing retrieval in isolation is insufficient. Even a perfect retriever, measured by traditional IR metrics like NDCG or MRR, may produce results that are poorly suited for LLM reasoning. This insight catalyzed a wave of research focused on training bridge tools: rerankers, query reformulators, and document selectors that explicitly align retrieval outputs with LLM preferences.
The Architecture of Preference Translation
BGM [159] crystallizes this new paradigm. The core observation is stark: there exists a systematic “preference gap” between what traditional retrievers optimize for (surface-level relevance, lexical overlap) and what LLMs find useful for reasoning (contextual coherence, inferential support). BGM addresses this by training a T5-XXL “bridge model” that sits between a frozen retriever and a frozen generator (PaLM2-S), transforming the retriever’s output into an LLM-friendly context. The training methodology is sophisticated: Stage 1 uses supervised learning on synthetically generated preference data (documents that improve LLM task performance vs. those that don’t), while Stage 2 employs reinforcement learning where the frozen LLM’s final task success provides the ultimate reward signal. This two-stage approach allows the bridge model to first learn coarse-grained preferences efficiently, then fine-tune its policy for actual downstream impact. The results are impressive: on HotpotQA, the bridged system achieves 35.6% exact-match accuracy compared to 25.8% with the best prior retriever—a relative improvement of 38%. What makes BGM architecturally significant is its demonstration that specialized adaptation layers can be more effective than end-to-end fine-tuning. Rather than trying to make a single retriever simultaneously satisfy IR metrics and LLM preferences (which may be in tension), BGM decomposes the problem: the retriever handles broad recall, while the bridge model handles preference alignment. This modularity is a recurring theme in the most successful T2 systems.
Synthesis: The Multi-Tool Ecosystem
@@
Collectively, these advances reveal a new architectural pattern: cascaded tool adaptation. Rather than a single monolithic retriever, state-of-the-art T2 systems now employ a pipeline of specialized tools: query reformulators, retrievers, selectors, each trained using different aspects of the frozen LLM’s behavior as supervision. This decomposition offers multiple advantages:
@@
Yet this raises a fundamental question: if tools can learn from tools, which learn from frozen LLMs, how deep can this hierarchy go before compounding errors overwhelm the benefits? This question remains open, though empirical results suggest that 2-3 stages of tool adaptation (e.g., query reformulator $\rightarrow$ retriever $\rightarrow$ reranker) strikes a good balance.

### v3

@@ 5 Tool Adaptation
Tool adaptation shifts the optimization target from the agent itself to its ecosystem. Instead of modifying the agent’s parameters through fine-tuning or reinforcement learning, this paradigm targets the external components (pre-trained models, retrievers, planners, or executors) that the agent invokes through language or code. Methods in this category (1) employ pre-trained machine learning models as plug-and-play components, ranging from simple classifiers to complex LLM-based subagents as discussed in §4; or (2) use the agent’s outputs as supervision or reinforcement signals to train, align, or refine the tool itself.
@@
where $\mathcal{O}_{\text{tool}}$ quantifies task-specific or environment-driven improvements that are independent of the agent, such as retrieval accuracy or planning efficiency, while $\mathcal{O}_{\text{agent}}$ incorporates agent-derived supervision, where the agent’s outputs provide learning signals to refine or align the tool. Here, $\mathcal{T}^{*}$ denotes the optimized tool configuration that maximizes the respective objective, illustrating how tool adaptation complements agent-level optimization within the broader agent-tool ecosystem.
@@ 5.1 T1: Agent-Agnostic Tool Adaptation
The foundational architecture for tool-augmented systems uses pre-trained models as plug-and-play tools for frozen agents. The agent orchestrates tool usage through prompting alone, never updating its parameters, while relying on tools trained independently on diverse data sources before deployment.
[…]
@@ 5.1.1 Foundational Systems and Architectures
Early systems established the architectural foundations for how frozen agents orchestrate external models. These works illustrate distinct mechanisms (functional, prompt-based, code-based, and graph-based) that shaped the design space for modern tool-augmented AI systems.
@@
Operator-Learning Tools: Neural Operators [125] (JMLR). Before large-scale LLM-based orchestration emerged, Neural Operators represented an early example of agent-agnostic tool learning: models trained to approximate mappings between infinite-dimensional function spaces, serving as differentiable surrogates for complex simulators. Unlike conventional neural networks tied to discrete grids, Neural Operators are discretization-invariant: they learn the underlying operator itself, not its finite discretization, and can generalize across resolutions and geometries. The Fourier Neural Operator (FNO) achieves $\mathcal{O}(J\log J)$ inference via spectral convolution and outperforms classical solvers on Navier-Stokes, Darcy flow, and elasticity equations by orders of magnitude in speed. FNO and its variants (Graph-, Low-rank-, Multipole-NO) represent the first wave of “frozen tools” that agents can query repeatedly for reasoning, planning, or control without retraining. In modern agentic pipelines, they serve as plug-in surrogates: fast, differentiable black-box functions invoked within decision or inference loops.
@@
HuggingGPT [126] (NeurIPS 2023) introduced the orchestration paradigm by allowing ChatGPT to command 1000+ machine learning models from HuggingFace Hub without any fine-tuning. The frozen LLM executes a four-stage workflow: task planning (decomposing user requests), model selection (choosing from tool descriptions), task execution (invoking models), and response generation (synthesizing outputs). The architecture shows that tool descriptions in natural language suffice for the frozen agent to coordinate complex multimodal workflows. On composite cross-modal tasks, HuggingGPT demonstrates that orchestrating specialized vision, speech, and language models can close the gap between GPT-3.5 and much larger multimodal models. The primary limitation lies in latency from sequential LLM calls and token length constraints for tool descriptions. ViperGPT [127] (ICCV 2023) introduced code generation as the orchestration mechanism. The frozen GPT-3 Codex generates Python code that composes vision models (GLIP for detection, SAM for segmentation, MiDaS for depth estimation) into executable programs. The code-based approach achieves strong zero-shot performance on GQA visual reasoning, outperforming end-to-end models by 10–15% on compositional tasks. Python functions provide more flexible tool composition than fixed API calls. Each tool exposes simple functions like find(image, object_name) or compute_depth(image), which Codex chains programmatically without learning tool-specific interfaces. SciToolAgent [128] (Nature Computational Science 2025) scales tool orchestration to scientific domains through graph-based organization. The frozen GPT-4o accesses 500+ biology, chemistry, and materials science tools via SciToolKG, a knowledge graph encoding tool metadata, dependencies, and safety constraints. Graph-based retrieval for tool selection achieves 94% accuracy on scientific query benchmarks, representing a 15–20% improvement over GPT-4o without tool access. The system successfully automates protein engineering workflows chaining ESMFold for structure prediction, BLAST for sequence alignment, and custom analysis tools. Structured knowledge graphs address the scalability challenges inherent in prompt-based descriptions.
@@
These foundational systems illustrate the dominant integration patterns. HuggingGPT exemplifies prompt-based orchestration, where the agent parses tool calls from text. ViperGPT uses code generation, exposing tools as Python functions. SciToolAgent uses knowledge graph retrieval, using RAG to select from structured tool graphs. A fourth common pattern is multimodal bridging, which converts non-textual modalities into text representations; for example, Visual ChatGPT’s [129] prompt manager serializes vision operations as text API calls. The usability of these patterns depends on clear interface design, such as programmatic function signatures (e.g., find_object(image: PIL.Image, ...)), structured JSON schemas, or simple natural language descriptions. Execution modes are similarly varied, ranging from direct API calls and code generation to HTTP requests and command-line invocations.
@@
Model Context Protocol (MCP) and Code Execution Environments [130]. As large-scale agent ecosystems began connecting thousands of heterogeneous tools, the Model Context Protocol (MCP) emerged as an open standard for unifying how agents interface with external systems. MCP provides a universal API layer that allows frozen agents to discover, invoke, and coordinate tools across domains using a consistent schema, rather than embedding long tool definitions directly into the model’s context. Anthropic’s “Code Execution with MCP” paradigm introduced an execution-centric design in which the agent writes executable code to interact with MCP servers instead of performing token-level tool calls. Agents can thereby load only the necessary tool definitions, filter or aggregate data within a sandboxed environment, and pass compact results back to the model, substantially reducing context usage while preserving compositionality. MCP represents a scalable T1-style tool adaptation infrastructure that decouples execution from inference, while the code-execution mode bridges toward T2-style optimization by dynamically improving efficiency under frozen agents.
@@ 5.1.2 Categories and Training Methodologies
While the preceding systems define how agents invoke tools, the choice of which tool modalities to deploy shapes the entire adaptation surface.
Tool adaptation encompasses a range of pre-trained model categories. We organize these by modality and highlight how their training methodology determines their amenability to T2 optimization: models producing structured, evaluable outputs (code, retrieval results) are most naturally upgraded to T2, because the frozen agent can score their outputs automatically; models producing perceptual features (vision, speech) are harder to supervise via T2 and remain predominantly T1 plug-ins.
@@
Vision models dominate T1 deployments as plug-and-play tools. CLIP [131], trained contrastively on 400M image-text pairs, provides zero-shot classification and semantic understanding through frozen encoders. SAM [132], trained on 11M images with 1B masks via human-in-the-loop data engines, provides promptable segmentation with point, box, or mask inputs. SAM-CLIP [133] merges these capabilities through multi-task distillation with frozen teachers, achieving +6.8% mIoU on Pascal VOC for zero-shot semantic segmentation while retaining both parent models’ strengths. These vision tools require no task-specific fine-tuning - frozen agents invoke them directly via APIs for image understanding, segmentation, and classification tasks.
@@
Speech and audio tools rely on massive pre-training for robust performance. Whisper [134], trained on 680K hours of multilingual audio with weak supervision, provides speech recognition, translation, and language identification as a frozen API for multimodal agents. The encoder-decoder Transformer architecture supports zero-shot transcription across languages and domains, with strong robustness to accents, noise, and technical terminology. Agents simply pass audio inputs to the frozen Whisper model and process text outputs without any model adaptation.
@@
Code execution tools encompass models that learn to compose and execute functions through code. CodeAct [58] shows that representing tool use in executable Python rather than static JSON improves compositional reasoning, achieving over 20% higher success rates on API-Bank benchmarks. The dynamic nature of code allows agents to flexibly construct, parameterize, and combine tools without predefined schemas.
@@
Search and retrieval tools comprise pre-trained dense retrievers such as DPR [135], ColBERT [136], Contriever [137], and e5 [138], often deployed as frozen components within retrieval-augmented generation pipelines. These bi-encoder models, trained on passage ranking tasks, support semantic search over large corpora.
@@
Scientific tools extend capabilities to specialized fields. AlphaFold2 [139] and ESMFold [140] provide protein structure prediction from sequences. Materials science models like CGCNN [141] predict crystal properties. Molecule representation learning approaches [142, 143, 144, 145, 146, 147, 148] have been developed for molecular property prediction, whereas some encoder-decoder frameworks [149, 150] aim to predict transcriptional profiles elicited by chemical perturbations. These tools represent years of domain-specific model development, deployed as-is for frozen agents tackling scientific queries.
@@ 5.2 T2: Agent-Supervised Tool Adaptation
The T2 paradigm inverts the standard adaptation question: rather than asking “how can we modify the agent to better use its tools?” (the A1/A2 question), T2 asks “how can we modify the tools to better serve a fixed agent?” This reframes the expensive, monolithic foundation model as a stable source of supervision rather than the target of optimization.
Training or fine-tuning billion-parameter foundation models is computationally expensive and risks catastrophic forgetting. Peripheral tools (retrievers, planners, memory modules) are orders of magnitude smaller and cheaper to train. T2 exploits this asymmetry: the frozen agent provides supervision signals derived from its pre-trained knowledge, while the tools learn to reshape information into the form the agent can best exploit.
The evolution of T2 methods from 2023 to 2025 shows a progression from using internal proxy signals (perplexity, preferences) to train passive retrieval tools, to using verifiable outcome signals (task success, accuracy gains) to train active, multi-turn agentic tools.
@@ 5.2.1 Earlier Methods: From Proxy Signals to Structured Preferences
The earliest T2 methods emerged from the RAG community, where researchers sought to optimize dense retrievers for compatibility with large language models. These works established the principle that a frozen LM’s internal computations could serve as supervision, but they also reveal the limitations of relying on proxy metrics that may not align with downstream task objectives.
REPLUG [11] (NAACL 2024) introduced a general framework for adapting frozen language models through black-box supervision, using perplexity reduction as a training signal for the retriever. If conditioning the LM on a retrieved document lowers its perplexity for a given query, the document likely provides informative context. Formally, the retriever is optimized to align its retrieval distribution with the distribution induced by the LM’s perplexity-based preferences:
@@
where $P_{\text{LM-perplexity}}(d|q)$ reflects how strongly each document reduces the LM’s perplexity when conditioned on query $q$. REPLUG thus enabled retrieval adaptation without parameter access to the LM, establishing a family of agent-supervised methods that optimize external modules based solely on frozen-agent feedback. BLADE [151] (AAAI 2025) further extended this paradigm by replacing traditional retrievers with domain-specific models that synthesize auxiliary knowledge: it couples a frozen general LLM with a small, domain-specific LM optimized via Bayesian Prompted Optimization (BPO). The small LM learns to generate domain-relevant knowledge and soft prompts that improve the black-box LLM’s responses, extending REPLUG’s black-box adaptation principle from retrieval to generative, domain-specialized tool co-adaptation. BBox-Adapter [152] (ICML 2024) reframed adaptation as an energy-based modeling problem, introducing a ranking-based noise-contrastive estimation loss and an online update framework to align outputs from black-box APIs like GPT-3.5 without access to internal probabilities. These methods collectively shifted the focus from likelihood-based alignment to utility-driven adaptation, anticipating the reinforcement-learning-based search and reasoning frameworks that followed. proxy-tuning [153] (COLM 2024) extends black-box, agent-supervised adaptation to decoding time: a small tuned “expert” and its untuned “anti-expert” provide logit offsets that steer a frozen large LM without modifying its weights, effectively training a lightweight steering tool under a frozen agent (T2). EVOR [154] (EMNLP 2024) further extends to the domain of code generation, formulating retrieval and knowledge evolution as a co-adaptive process driven by execution feedback from a frozen LLM.
@@
While black-box adaptation methods such as REPLUG and BBox-Adapter relied on indirect proxy signals like perplexity or ranking scores, subsequent studies [29, 155, 156, 157, 158] moved toward explicit preference-based supervision that better reflects task utility. AAR [29] (ACL 2023) introduced augmentation-aware retrieval, where a frozen source LM (Flan-T5-250M) constructs preference pairs by comparing documents that most improve its own likelihood against human-annotated references. The retriever is then trained to reproduce these preferences via a contrastive loss, yielding a signal that directly encodes the LM’s notion of “helpful context.” These LM-derived preferences transfer effectively across architectures and scales, improving even 175B-parameter models. RA-DIT [155] (ICLR 2023) formalized this idea by defining document utility as the log-probability gain for producing correct answers:
@@
training retrievers to prefer documents that yield higher expected gains. The preference-based approach aligns retrieval more directly with downstream reasoning objectives while still operating through a single forward pass of the frozen LM. Together, these works mark a conceptual shift from proxy-based to task-aligned supervision, setting the stage for reinforcement-style feedback and multi-turn optimization in later frameworks.
Multi-stage architectures: distilling complex preferences.
@@
LLM-R [24] (EACL 2024) introduced a multi-stage distillation pipeline that increases the fidelity of training signals. Rather than training the retriever directly on the frozen LM’s outputs, LLM-R first trains an intermediate cross-encoder “reward model” to capture the frozen LM’s nuanced preferences over in-context examples. The reward model, which can afford to be slow and expressive because it is used only during training, is then distilled into a fast bi-encoder retriever. The key insight is that the complexity of the training signal need not be constrained by inference-time efficiency requirements. By decoupling the preference modeling from the final retrieval tool, LLM-R achieves both high-quality supervision and fast deployment.
UPRISE [159] (EMNLP 2023) extends this paradigm beyond documents to prompts, training a prompt retriever using the frozen LLM’s task performance across diverse tasks. By training on multiple tasks simultaneously, UPRISE learns a generalizable meta-skill: selecting prompts that improve LLM performance in zero-shot settings. The cross-task transfer (+8.5% on reading comprehension, +14.6% on paraphrase detection) suggests that T2-trained tools can internalize abstract principles of “what helps an LM” rather than memorizing task-specific heuristics.
@@
By 2024, a consensus emerged that optimizing retrieval in isolation is insufficient. Even a retriever that scores well on traditional IR metrics (NDCG, MRR) may produce results poorly suited for LLM reasoning. The observation motivated research on bridge tools: rerankers, query reformulators, and document selectors that align retrieval outputs with LLM preferences.
The architecture of preference translation.
BGM [160] addresses the systematic “preference gap” between what traditional retrievers optimize for (surface-level relevance, lexical overlap) and what LLMs find useful for reasoning (contextual coherence, inferential support). BGM addresses this by training a T5-XXL “bridge model” that sits between a frozen retriever and a frozen generator (PaLM2-S), transforming the retriever’s output into an LLM-friendly context.
Stage 1 uses supervised learning on synthetically generated preference data (documents that improve LLM task performance vs. those that do not), while Stage 2 employs reinforcement learning where the frozen LLM’s final task success provides the reward signal. On HotpotQA, the bridged system achieves 35.6% exact-match accuracy compared to 25.8% with the best prior retriever, a relative improvement of 38%.
BGM shows that specialized adaptation layers can be more effective than end-to-end fine-tuning. Rather than making a single retriever simultaneously satisfy IR metrics and LLM preferences, BGM decomposes the problem: the retriever handles broad recall, while the bridge model handles preference alignment. The same modularity recurs in the most successful T2 systems.
Synthesis: the multi-tool ecosystem.
@@
These advances reveal an architectural pattern: cascaded tool adaptation. State-of-the-art T2 systems now employ a pipeline of specialized tools (query reformulators, retrievers, selectors), each trained using different aspects of the frozen LLM’s behavior as supervision. The decomposition offers several advantages:
@@
An open question is how deep this hierarchy can go before compounding errors overwhelm the benefits. Empirical results suggest that 2–3 stages of tool adaptation (e.g., query reformulator $\rightarrow$ retriever $\rightarrow$ reranker) strike a good balance.